OpenStudy (anonymous):

lim 0 goes to theta of sin(3 theta)/(theta)

OpenStudy (luigi0210):

Welcome to Openstudy! This is the question right? \[\lim_{x \rightarrow 0} \frac{\sin \theta}{\theta}\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

lim theta goes to 0

OpenStudy (luigi0210):

\[\lim_{\theta \rightarrow 0} \frac{\sin 3 \theta}{\theta}\]

OpenStudy (anonymous):

Yeah that's the question

OpenStudy (anonymous):

theta is measured in radians

OpenStudy (luigi0210):

@SithsAndGiggles This is simple for you

OpenStudy (anonymous):

I must estimate the limit

OpenStudy (anonymous):

Recall that \[\lim_{\theta\to0}\frac{\sin a\theta}{a\theta}=1\] How can you get a 3 in the denominator without fundamentally altering the function?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

Well if you were to multiply by 1, you'd have the same thing right? Fortunately for use, \(\dfrac{3}{3}=1\), so \[\frac{\sin3\theta}{\theta}=\frac{\sin3\theta}{\theta}\cdot\frac{3}{3}\]

OpenStudy (anonymous):

I have that...not sure where to go from there

OpenStudy (anonymous):

Then you also have \[3\lim_{\theta\to0}\frac{\sin3\theta}{3\theta}\]

OpenStudy (anonymous):

What's the next step?

OpenStudy (anonymous):

See the previous comment: \(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles Recall that \[\lim_{\theta\to0}\frac{\sin a\theta}{a\theta}=1\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Got it, the answer is 3